# Droplet coalescence kinetics: Coalescence mechanisms and thermodynamic non-equilibrium effects under isothermal and non-isothermal conditions

### v1

@@ III Numerical simulations and analysis
Two initial static droplets with equal sizes is set as:
$\begin{array}[]{l}\rho(x,y)={\rho_{\rm L}}-\frac{{({\rho_{\rm L}}-{\rho_{\rm V}})}}{2}\tanh\left[{\frac{{\sqrt{{{(x-{x_{0\rm l}})}^{2}}+{{(y-{y_{0\rm l}})}^{2}}}-{r_{0}}}}{{0.5w}}}\right]\\ \mathrm{\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ }-\frac{{({\rho_{\rm L}}-{\rho_{\rm V}})}}{2}\tanh\left[{\frac{{\sqrt{{{(x-{x_{0\rm r}})}^{2}}+{{(y-{y_{0\rm r}})}^{2}}}-{r_{0\rm r}}}}{{0.5w}}}\right].\end{array}$ (8)
@@ III.1 Differences in HNE effects and coalescence mechanisms
Figure 1: Density distribution at eight identical time instances under isothermal and non-isothermal conditions: (A) Isothermal case; (B) Non-isothermal case. Each panel corresponds to a partial simulation area with ${N_{x}}\times{N_{y}}=256\times 256$.
@@
Figure 1 shows the spatial distribution of density at eight characteristic moments with similar morphology under isothermal and non-isothermal conditions. The droplet coalescence process is significantly slower in the non-isothermal case compared to the isothermal case. Specifically, under non-isothermal conditions, the initiation of the coalescence process is delayed, cut-through occurs more slowly, the rapid coalescence phase lasts longer, and exhibits noticeable damped oscillations.
(i) Slower initiation and cut-through processes: The coalescence initiation duration under non-isothermal conditions is much longer than that under isothermal conditions. For the isothermal case, the cut-through time of two static droplets is $t_{c}=1.8$ [see Fig. 1 Ab], while for the non-isothermal case, it is $t_{c}=5.6$ [see Fig. 1 Bb]. The time difference in both cases is $\Delta t_{c}=3.8$. The differences in cut-through time can be more obviously reflected in the evolution of morphological factor. As shown in Fig. 2, in both cases, at $t=t_{c}$, the morphological factor $L_{d}=L_{y}/L_{x}$ and $L_{y}$ begin to increase, where $L_{x}$ and $L_{y}$ are the long and short semi-axes of the new droplet, respectively.
(ii) Longer duration of the rapid coalescence process: As shown in Fig. 1 and Fig. 2, under isothermal conditions, the time at which the new droplet first forms an unstable spherical shape, corresponding to the morphological factor $L_{d}$ first equaling 1, is $t_{l1}=7.0$. Then the time difference between cut-through and the first formation of an unstable spherical droplet is $\Delta t_{l1c}=t_{l1}-t_{c}=5.2$. Under non-isothermal conditions, the time for the new droplet to first forms an unstable spherical shape is $t_{l1}=15.2$, and the time difference between contact cut-through and the first formation of an unstable spherical droplet is $\Delta t_{l1c}=t_{l1}-t_{c}=9.6$. The time difference for the first formation of an unstable spherical droplet in both cases is $\Delta t_{l1}=8.8$. Under non-isothermal conditions, the growth rate of the morphological factor $L_{d}$ of the new droplet is smaller than under isothermal conditions, because the slopes of both the long semi-axis $L_{x}$ and the short semi-axis $L_{y}$ of the droplet are smaller under non-isothermal conditions [see Fig. 2].
(iii) Greater oscillation amplitude and longer oscillation period: As shown in Fig. 2(a), under isothermal conditions, the time at which $L_{d}$ reaches its maximum is $t_{l\rm max}=9.6$, whereas under non-isothermal conditions, $t_{l\rm max}=20.0$, with a difference of $\Delta t_{l\rm max}=10.4$. After the first oscillation period, under isothermal conditions, the oscillation amplitude of $L_{x}$ and $L_{y}$ is almost zero, as shown in Fig. 2(b). In contrast, under non-isothermal conditions, small oscillations persist. Furthermore, the maximum value of $L_{d}$ under non-isothermal conditions is greater than that under isothermal conditions. This can be attributed to the more complex stress and flow field distributions near the vapor-liquid interface under non-isothermal conditions, which prevent the liquid-vapor interface from stabilizing easilyChen, Hong, and Zang (2024).
@@
To explain the above differences, Fig. 3 presents the spatial distribution along the horizontal midline of various quantities: temperature $T$, latent heat of phase change $L=\Delta H=H_{t}-H_{0}$ (with enthalpy $H=e_{T}+P/\rho$), velocity in the coalescence direction $u_{x}$, density $\rho$, pressure $P$, the $xx$-component of the total pressure tensor $\bm{\Gamma}=P\mathbf{I}+\bm{\Lambda}$ (denoted as $\Gamma_{xx}$), and the $xx$-component of the pressure tensor contribution from the macroscopic gradient $\bm{\Lambda}$ (denoted as $\Lambda_{xx}$). In the isothermal case, it is assumed that the system is in contact with a sufficiently large heat bath, resulting in instantaneous heat exchange with the environmentGan et al. (2011). This assumption leads to no temperature gradient within the system. In contrast, the non-isothermal case incorporates the temperature field’s inhomogeneity, introducing additional HNE and TNE effects, such as the temperature-induced Marangoni effect and heat conduction. These effects significantly alter the system’s dynamics and flow field morphology.
@@
Figure 5: Time evolution of the average coalescence velocity $\bar{u}$, average coalescence acceleration $\bar{a}$ and $dL/dt$, under both isothermal and non-isothermal conditions. Here $\bar{u}={\sum{\rho(x,y,t){{u}_{x}}(x,y,t)}}/{\sum{\rho(x,y,t)}}$ and $\bar{a}=d\bar{u}/dt$.
@@
(i) Under isothermal conditions, when $t\leq 3.8$, the velocity in the vapor phase along the vertical direction of coalescence (the $y$-direction) converges toward the saddle point in the saddle region between the two droplets. No vortices are formed at the liquid-vapor interface in this region. In contrast, under non-isothermal conditions, when $t\leq 10.8$, the velocity in the vapor phase outside the saddle point diverges along the vertical direction of coalescence. This divergence creates vortices with opposite rotational directions at the liquid-vapor interface in the saddle region. The non-isothermal case results in a complex flow pattern near the interface, with the Marangoni effect significantly enhancing the flow outside the interface.
(ii) Under isothermal conditions, due to the faster release of surface energy, the peak of the velocity in merging direction is higher compared to the non-isothermal case. As shown in Fig. 5, the coalescence speed under isothermal conditions starts faster and reaches its peak earlier. Specifically, the time when the coalescence speed reaches its maximum for the isothermal case is $t_{\rm umax}=3.8$, while for the non-isothermal case, $t_{\rm umax}=10.8$. Furthermore, the peak value $\bar{u}_{\rm max}$ is larger in the isothermal case. However, under isothermal conditions, the coalescence speed decays more rapidly, particularly during the damping oscillation phase ($t\geq t_{\rm l1}$). After the first oscillation cycle, the coalescence speed under isothermal conditions almost decays to zero, while under non-isothermal conditions, the coalescence speed continues to oscillate for several cycles before gradually decaying to zero.
@@ III.2 Differences in TNE effects
Figure 6: Spatiotemporal evolution characteristics of the system’s NOMF: (A) Isothermal conditions, (B) Non-isothermal conditions. Each sub-figure represents a partial simulated area with ${N_{x}}\times{N_{y}}=256\times 256$.
@@
The key differences between the two cases are as follows: (i) In the isothermal case, the spatial distributions of $\Delta^{\ast}_{2xx}$ and $\Delta^{\ast}_{2yy}$ are not completely antisymmetric due to the different directions of the velocity vortices. In contrast, these distributions are fully antisymmetric in the non-isothermal case. (ii) In the non-isothermal case, the spatial distribution of $\Delta^{\ast}_{2xy}$ is more complex due to the increased variety of non-equilibrium driving forces. $\Delta^{\ast}_{2xy}$ in the non-isothermal case evolves from an interior-exterior antisymmetric double quadrupole, while in the isothermal case, it evolves from an antisymmetric single quadrupole structure. The coupling of multiple TNE effects in the non-isothermal system makes the spatiotemporal evolution of non-equilibrium behaviors more complex.
Figure 7 illustrates the time evolution of several non-equilibrium intensities, including the total non-equilibrium intensity $\bar{D}^{\ast}$, NOMF intensity $\bar{D}^{\ast}_{2}$, NOEF intensity $\bar{D}^{\ast}_{3,1}$, strength of flux of NOMF $\bar{D}^{\ast}_{3}$, and strength of flux of NOEF $\bar{D}^{\ast}_{4,2}$ in both isothermal and non-isothermal systems. The definitions of these intensities are based on the sums of the magnitudes of the respective non-equilibrium flux components, weighted by the density of the system: $\bar{D}_{m,n}^{*}={\sum{\rho\left|\bm{\Delta}_{m,n}^{*}\right|}}/{\sum{\rho}}$, and ${{\bar{D}}^{*}}={\sum\rho\sqrt{\bm{\Delta}_{2}^{*2}+\bm{\Delta}_{3}^{*2}+\bm{\Delta}_{3,1}^{*2}+\text{ }\bm{\Delta}_{4,2}^{*2}}}/{\sum\rho}$.
The main differences between the two cases are as follows:
@@
This is because, in the early stage of the relaxation process, when the droplets approach each other, the velocity gradient is negligible, and the temperature gradient due to latent heat of phase change governs the system’s evolution. After the droplets make contact, the relative velocity between particles increases, causing the velocity gradient to dominate. As shown in Fig. 5, at $t=5.0$, the droplets begin to make contact but have not yet merged, marking the moment when the boundary length rapidly decreases. In the later stages of coalescence, $\bar{D}^{\ast}_{3,1}$ and $\bar{D}^{\ast}_{3}$ decay smoothly to zero, while $\bar{D}^{\ast}_{2}$ and $\bar{D}^{\ast}_{4,2}$ exhibit damped oscillations and also decay to zero as the droplet shape evolves. In summary, the increasing number of driving factors in the non-isothermal case adds the interplay and complexity to the system, as evidenced by the alternating dominance of behaviors driven by temperature gradients ($\bm{\Delta}^{\ast}_{3}$, $\bm{\Delta}^{\ast}_{3,1}$) and velocity gradients ($\bm{\Delta}^{\ast}_{2}$, $\bm{\Delta}^{\ast}_{4,2}$).
(v) Higher-order TNE effects in isothermal case: In the isothermal case, despite no temperature gradient, $\bar{D}^{\ast}_{3,1}$ and $\bar{D}^{\ast}_{3}$ are not exactly zero due to higher-order effects, such as velocity gradients and density gradients, which also contribute to the heat flux intensityGan et al. (2022).
@@ III.3 Differences in entropy production mechanisms
Figure 9: Time evolution of the entropy production rate under both isothermal and non-isothermal conditions.
@@ III.4 Effects of initial distance on the coalescence process
Figure 12 shows how the initial distance $r_{d}$ between two droplets affects the cut-through time $t_{c}$. In both isothermal and non-isothermal cases, the cut-through time $t_{c}$ increases with the initial droplet distance $r_{d}$, following a power-law relationship. For identical parameters, the exponent in the non-isothermal case is smaller than in the isothermal case. This behavior is due to the similarity in the system’s dynamic. Because, in the non-isothermal case, the temperature gradient induces energy dissipation mechanisms, such as the Marangoni effect and heat conduction, which delay the droplet coalescence stage.
@@
The power exponent decreases as the surface tension coefficient $K$ increases. In other words, a larger surface tension coefficient $K$ results in a smaller crossing time $t_{c}$. As shown in Fig. 13, $t_{c}$ and $K$ follow a negative power function relationship. This can be attributed to two factors: First, surface tension is the primary driving force for droplet coalescence, regardless of whether the case is isothermal or non-isothermal. A larger surface tension coefficient $K$ increases the surface tension, accelerating droplet approach and shortening the crossing time. Second, a larger surface tension coefficient increases the liquid-vapor interface widthWagner and Pooley (2007) $w=\sqrt{{2K}/{({{{T}_{c}}}/{T}-1)}}$, effectively reducing $r_{d}$ and shortening the crossing time.
@@
${{t}_{c}}=2.68\times{{10}^{-21}}{{K}^{-5.584}}\quad(r_{d}=15\Delta x).$
@@
Figure 14: Effects of $r_{d}$ on the total non-equilibrium strength $\bar{D}^{\ast}$ with varying under both isothermal and non-isothermal conditions.

### v3

@@ III Simulations and results
Two initially stationary droplets of identical size are positioned as:
$\begin{array}[]{l}\rho(x,y)={\rho_{\rm L}}-\frac{{({\rho_{\rm L}}-{\rho_{\rm V}})}}{2}\tanh\left[{\frac{{\sqrt{{{(x-{x_{0\rm l}})}^{2}}+{{(y-{y_{0\rm l}})}^{2}}}-{r_{0}}}}{{0.5w}}}\right]\\ \mathrm{\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ }-\frac{{({\rho_{\rm L}}-{\rho_{\rm V}})}}{2}\tanh\left[{\frac{{\sqrt{{{(x-{x_{0\rm r}})}^{2}}+{{(y-{y_{0\rm r}})}^{2}}}-{r_{0}}}}{{0.5w}}}\right].\end{array}$ (14)
@@ III.1 Differences in HNE effects and coalescence mechanisms
Figure 1: Density distribution at eight characteristic moments with similar morphology under isothermal and non-isothermal conditions: (A) Isothermal case; (B) Non-isothermal case. Here, each sub-figure represents a partial simulated area with grid dimensions ${N_{x}}\times{N_{y}}=256\times 256$.
Figure 1 shows the spatial distribution of density at eight characteristic moments with similar morphology under isothermal and non-isothermal conditions. The droplet coalescence process is significantly slower in the non-isothermal case compared to the isothermal case. Specifically, under non-isothermal condition, the initiation of the coalescence process is delayed, cut-through occurs more slowly, the rapid coalescence phase lasts longer, and exhibits noticeable damped oscillations.
@@
(i) Slower initiation and cut-through processes: The coalescence initiation duration under non-isothermal condition is much longer than that under isothermal condition. For the isothermal case, the cut-through time of two static droplets is $t_{c}=1.8$ [see Fig. 1 Ab], while for the non-isothermal case, it is $t_{c}=5.6$ [see Fig. 1 Bb]. The time difference in both cases is $\Delta t_{c}=3.8$. The differences in cut-through time can be more obviously reflected in the evolution of morphological factor. As shown in Fig. 2, in both cases, at $t=t_{c}$, the morphological factor $L_{d}=L_{y}/L_{x}$ and $L_{y}$ begin to increase, where $L_{x}$ and $L_{y}$ are the long and short semi-axes of the new droplet, respectively.
(ii) Longer duration of the rapid coalescence process: As shown in Fig. 1 and Fig. 2, under isothermal condition, the time at which the new droplet first forms an unstable spherical shape, corresponding to the morphological factor $L_{d}$ first equaling 1, is $t_{l1}=7.0$. Then the time difference between cut-through and the first formation of an unstable spherical droplet is $\Delta t_{l1c}=t_{l1}-t_{c}=5.2$. Under non-isothermal condition, the time for the new droplet to first forms an unstable spherical shape is $t_{l1}=15.2$, and the time difference between cut-through and the first formation of an unstable spherical droplet is $\Delta t_{l1c}=t_{l1}-t_{c}=9.6$. The time difference for the first formation of an unstable spherical droplet in both cases is $\Delta t_{l1}=8.2$. Under non-isothermal condition, the growth rate of the morphological factor $L_{d}$ of the new droplet is smaller than under isothermal condition, because the slopes of both the long semi-axis $L_{x}$ and the short semi-axis $L_{y}$ of the droplet are smaller under non-isothermal condition [see Fig. 2].
(iii) Greater oscillation amplitude and longer oscillation period: After coalescence, the newly formed larger droplet often exhibits a highly non-circular shape. Since the system is not initially at its minimum surface energy, surface tension drives the droplet toward restoring a circular configuration. During this relaxation process, due to fluid inertia, the droplet typically does not reach equilibrium in a single step. Instead, it undergoes damped oscillations around the spherical shape until its kinetic energy is fully dissipated. The timescale of these oscillations is governed jointly by viscous dissipation, heat transfer, and surface tension effects Qian and Law (1997). As shown in Fig. 2(a), under isothermal condition, the time at which $L_{d}$ reaches its maximum is $t_{l\rm max}=9.6$, whereas under non-isothermal condition, $t_{l\rm max}=20.0$, with a difference of $\Delta t_{l\rm max}=10.4$. After the first oscillation period, under isothermal condition, the oscillation amplitude of $L_{x}$ and $L_{y}$ is almost zero, as shown in Fig. 2(b). In contrast, under non-isothermal condition, small oscillations persist. Furthermore, the maximum value of $L_{d}$ under non-isothermal condition is greater than that under isothermal condition. This can be attributed to the more complex stress and flow field distributions near the vapor-liquid interface under non-isothermal condition, which prevent the liquid-vapor interface from stabilizing easily Chen, Hong, and Zang (2024).
@@
To explain the above differences, Fig. 3 presents the spatial distribution along the horizontal midline of various quantities: temperature $T$, latent heat of phase change $\Delta H=H_{t}-H_{0}$ (with enthalpy $H=\epsilon+P/\rho$, the internal energy density ${\epsilon}=\rho T-a{\rho^{2}}+{{K{{\left|{\bm{\nabla}\rho}\right|}^{2}}}/2}$, and $H_{t}$ and $H_{0}$ being the enthalpies at time $t$ and initial state), velocity in the coalescence direction $u_{x}$, density $\rho$, pressure $P$, the $xx$-component of the total pressure tensor $\bm{\Gamma}=P\mathbf{I}+\bm{\Lambda}$ (denoted as $\Gamma_{xx}$), and the $xx$-component of the pressure tensor contribution from the macroscopic gradient $\bm{\Lambda}$ (denoted as $\Lambda_{xx}$).
In the isothermal case, it is assumed that the system is in contact with a sufficiently large heat bath, resulting in instantaneous heat exchange with the environment Gan et al. (2011). This assumption leads to no temperature gradient within the system. In contrast, the non-isothermal case incorporates the temperature field’s inhomogeneity, introducing additional HNE and TNE effects, such as the temperature-induced Marangoni effect and heat conduction. These effects significantly alter the system’s dynamics and flow field morphology.
@@
Figure 5: Time evolution of the average coalescence velocity $\bar{u}$, average coalescence acceleration $\bar{a}$ and $dL/dt$, under both isothermal and non-isothermal conditions. Here, $\bar{u}$ is the statistical average of the velocity in the coalescencing direction within the left half region $\bar{u}={\sum{\rho(x,y,t){{u}_{x}}(x,y,t)}}/{\sum{\rho(x,y,t)}}$ and $\bar{a}=d\bar{u}/dt$.
@@
(i) Under isothermal condition, when $t\leq 3.8$, the velocity in the vapor phase along the vertical direction of coalescence (the $y$-direction) converges toward the saddle point in the saddle region between the two droplets. No vortices are formed at the liquid-vapor interface in this region. In contrast, under non-isothermal condition, when $t\leq 10.8$, the velocity in the vapor phase outside the saddle point diverges along the vertical direction of coalescence. This divergence creates vortices with opposite rotational directions at the liquid-vapor interface in the saddle region. The non-isothermal case results in a complex flow pattern near the interface, with the Marangoni effect significantly enhancing the flow outside the interface.
(ii) Under isothermal condition, due to the faster release of surface energy, the peak of the velocity in merging direction is higher compared to the non-isothermal case. As shown in Fig. 5, the coalescence speed under isothermal condition starts faster and reaches its peak earlier. Specifically, the time when the coalescence speed reaches its maximum for the isothermal case is $t_{u\rm max}=3.8$, while for the non-isothermal case, $t_{u\rm max}=10.8$. Furthermore, the peak value $\bar{u}_{\rm max}$ is larger in the isothermal case. However, under isothermal condition, the coalescence speed decays more rapidly, particularly during the damping oscillation phase ( $t\geq t_{l1}$). After the first oscillation cycle, the coalescence speed under isothermal condition almost decays to zero, while under non-isothermal condition, the coalescence speed continues to oscillate for several cycles before gradually decaying to zero.
@@ III.2 Differences in TNE effects
Figure 6: Spatiotemporal evolution characteristics of the system’s NOMF: (A) Isothermal condition, (B) Non-isothermal condition. Each sub-figure represents a partial simulated area with ${N_{x}}\times{N_{y}}=256\times 256$.
@@
The key differences between the two cases are as follows: (i) In the isothermal case, the spatial distributions of $\Delta^{\ast}_{2xx}$ and $\Delta^{\ast}_{2yy}$ are not completely opposite due to the different directions of the velocity vortices. In contrast, these distributions are fully antisymmetric in the non-isothermal case. (ii) In the non-isothermal case, the spatial distribution of $\Delta^{\ast}_{2xy}$ is more complex due to the increased variety of non-equilibrium driving forces. $\Delta^{\ast}_{2xy}$ in the non-isothermal case evolves from an interior-exterior antisymmetric double quadrupole, while in the isothermal case, it evolves from an antisymmetric single quadrupole structure. The coupling of multiple TNE effects in the non-isothermal system makes the spatiotemporal evolution of non-equilibrium behaviors more complex.
Figure 7 illustrates the time evolution of several non-equilibrium intensities, including the total non-equilibrium intensity $\bar{D}^{\ast}$, NOMF intensity $\bar{D}^{\ast}_{2}$, NOEF intensity $\bar{D}^{\ast}_{3,1}$, strength of flux of NOMF $\bar{D}^{\ast}_{3}$, and strength of flux of NOEF $\bar{D}^{\ast}_{4,2}$ in both isothermal and non-isothermal systems. The main differences between the two cases are as follows:
@@
This is because, in the early stage of the relaxation process, when the droplets approach each other, the velocity gradient is negligible, and the temperature gradient due to latent heat of phase change governs the system’s evolution. After the droplets make contact, the relative velocity between particles increases, causing the velocity gradient to dominate. As shown in Fig. 5, at $t=5.0$, the droplets begin to make contact but have not yet merged, marking the moment when the boundary length rapidly decreases. In the later stage of coalescence, $\bar{D}^{\ast}_{3,1}$ and $\bar{D}^{\ast}_{3}$ decay smoothly to zero, while $\bar{D}^{\ast}_{2}$ and $\bar{D}^{\ast}_{4,2}$ exhibit damped oscillations and also decay to zero as the droplet shape evolves. In summary, the increasing number of driving factors in the non-isothermal case adds the interplay and complexity to the system, as evidenced by the alternating dominance of behaviors driven by temperature gradients ($\bm{\Delta}^{\ast}_{3}$, $\bm{\Delta}^{\ast}_{3,1}$) and velocity gradients ($\bm{\Delta}^{\ast}_{2}$, $\bm{\Delta}^{\ast}_{4,2}$).
(v) Higher-order TNE effects in the isothermal case: Theoretically, the non-organized energy flux (NOEF) contains contributions from various orders of non-equilibrium effects, as expressed by $\bm{\Delta}_{3,1}^{*}=\bm{\Delta}_{3,1}^{*(1)}+\bm{\Delta}_{3,1}^{*(2)}+\cdots$, where $\bm{\Delta}_{3,1}^{*}=\bm{\Delta}_{3,1}^{*(1)}$ in the first-order model. According to theoretical derivation, $\bm{\Delta}_{3,1}^{*(1)}$ only includes contributions from the temperature gradient, whereas $\bm{\Delta}_{3,1}^{*(2)}$ incorporates both temperature and velocity gradient contributions Gan et al. (2022). Thus, in the isothermal case, despite the absence of a temperature gradient, $\bar{D}^{\ast}_{3,1}$ and $\bar{D}^{\ast}_{3}$ are not exactly zero. This phenomenon reflects the complex cross-coupling of multiple physical quantities in constitutive relations, which is one manifestation of the multiscale system’s complexity. This also indicates that the current DBM framework primarily focuses on first-order non-equilibrium effects, but naturally captures partial second-order non-equilibrium behaviors without requiring additional modifications.
Overall, as a phenomenon, non-equilibrium fundamentally embodies system complexity by revealing the nonlinear coupling among various physical mechanisms. As a driving effect, non-equilibrium significantly influences system evolution through its intrinsic, dominant, and integrative characteristics. Non-equilibrium quantities thus serve as sensitive probes for identifying competing mechanisms, providing earlier, more sensitive, and more precise detection of dominant processes and their transitions compared to conventional macroscopic variables.
@@ III.3 Differences in entropy production mechanisms
Figure 9: Temporal evolution of the entropy production rate under isothermal and non-isothermal conditions.
@@ III.4 Effects of initial distance on the coalescence process
Figure 12 shows how the initial distance $r_{d}$ between two droplets affects the cut-through time $t_{c}$. In both isothermal and non-isothermal cases, the cut-through time $t_{c}$ increases with the initial droplet distance $r_{d}$, following a power-law relationship. For identical parameters, the exponent in the non-isothermal case is smaller than that in the isothermal case. This behavior is due to the similarity in the system’s dynamic. Because, in the non-isothermal case, the temperature gradient induces energy dissipation mechanisms, such as the Marangoni effect and heat conduction, which delay the droplet coalescence stage.
@@
The power exponent decreases as the surface tension coefficient $K$ increases. In other words, a larger surface tension coefficient $K$ results in a smaller crossing time $t_{c}$. As shown in Fig. 13, $t_{c}$ and $K$ follow a negative power function relationship. This can be attributed to two factors: First, surface tension is the primary driving force for droplet coalescence, regardless of whether the case is isothermal or non-isothermal. A larger surface tension coefficient $K$ increases the surface tension, accelerating droplet approach and shortening the crossing time. Second, a larger surface tension coefficient increases the liquid-vapor interface width $w=\sqrt{{2K}/{({{{T}_{c}}}/{T}-1)}}$ Wagner and Pooley (2007), effectively reducing $r_{d}$ and shortening the crossing time.
@@
${{t}_{c}}=2.68\times{{10}^{-21}}{{K}^{-5.58}}\quad(r_{d}=15\Delta x).$
@@
Figure 14: Effects of $r_{d}$ on the total non-equilibrium strength $\bar{D}^{\ast}$ under both isothermal and non-isothermal conditions.